How do I solve ---- Find a number that if 16 is added to it, the ratio of the result to the original number is 4:3
Thank you!!
Cindy

Answer by Alan3354(69443)

How do I solve ---- Find a number that if 16 is added to it, the ratio of the result to the original number is 4:3
----------
n is "the number"
(n+16)/n = 4/3
3(n+16) = 4n
3n + 48 = 4n
n = 48